OpenStudy (anonymous):

Solve for the indicated variable. P = 2L + 2W; for L L=

OpenStudy (anonymous):

are there any answer choices?

OpenStudy (anonymous):

P = 2L + 2W 2W + 2L = P -2W -2W 2L=P-2W divide both sides by 2 L=(P-2W)/2

OpenStudy (anonymous):

@Annа

OpenStudy (anonymous):

I got the same answer as @larryboxaplenty